OpenStudy (anonymous):

PLEASE HELP! Find the general term of an arithmetic sequence that t6=15 and t11=-20. Hence, find the value of t17. No explanation is needed, just the working out. Will give medal :)

OpenStudy (academicgurusinc):

Use Use the general arithmetic sequence equation tn = t1 + (n-1)d -where where d is the common difference for eac succesive term of the sequence -where n describes the position of a term within a sequence In your case we know that t6=15 and t11 = -20 Therefore, t6=15=t1 + (6-1)d=t1 +5d ---> t1 = 15-5d t11=-20= t1 + (11-1)d=t1 +10d---> t1 = -20-10d Therefore, 15-5d=-20-10d--solve for d d=-7 Now you can solve for t1 by plugging d into either of the t6 or t11 equation: t11=-20=t1 + (10)(-7)--->t1 = 50 Now you can solve for t17 t17= t1 + (17-1)d= 50 + (16) (-7)=-62general arithmetic sequence equation tn = t1 + (n-1)d -where where d is the common difference for eac succesive term of the sequence -where n describes the position of a term within a sequence In your case we know that t6=15 and t11 = -20 Therefore, t6=15=t1 + (6-1)d=t1 +5d ---> t1 = 15-5d t11=-20= t1 + (11-1)d=t1 +10d---> t1 = -20-10d Therefore, 15-5d=-20-10d--solve for d d=-7 Now you can solve for t1 by plugging d into either of the t6 or t11 equation: t11=-20=t1 + (10)(-7)--->t1 = 50 Now you can solve for t17 t17= t1 + (17-1)d= 50 + (16) (-7)=-62

OpenStudy (academicgurusinc):

We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)